Sol Summary – February 20th

Sol: 10

Summary Title: More Puzzling Weather on Mars

Author’s name: Hannah Blackburn

Mission Status: Nominal

Sol Activity Summary: We woke up to an overcast sky and biscuits with sausage gravy. The crew worked on a jigsaw puzzle of Mars while drinking their morning coffee. The EVA 14 team left the Hab at 1300 and quickly realized that Spirit and Opportunity did not fully charge. Atila reset the RAM’s external breaker which solved the problem. Approximately one hour later, the EVA 14 team set out again towards Candor Chasma on Spirit and Curiosity. However, when they had nearly reached their destination, they noticed fast-moving clouds and decided to return to the Hab to prevent weather-related issues. Later in the afternoon, light precipitation occurred while the crew ate chocolate lava cake and continued to piece together the puzzle. We finished the evening by enjoying movies from the Hab collection. Apollo 13 is a crew favorite.

Look Ahead Plan: If weather and ground conditions allow, although it currently seems unlikely, part of the crew will go on EVA 15 to The Moon. We will continue to work on developing Standard Operating Procedures for future crews.

Anomalies in work: Spirit and Opportunity did not fully charge. Resetting the RAM’s external breaker fixed the issue.

Weather: Overcast, light flurries starting at 1530

Crew Physical Status: Nominal


Reports to be filed: Sol Summary, Journalist Report, GreenHab Report, Operations Report, EVA report, EVA request

Support Requested: We did not receive any cloth towels with the initial supplies and therefore are on our last roll of paper towels. We would appreciate more paper or cloth towels for cleaning.

Copyright © The Mars Society. All rights reserved. | Main Site